OpenStudy (anonymous):

How do you solve 3x^2=15?

OpenStudy (anonymous):

divide both sides by 3 x^2 = 5 x = +sqrt 5 , - sqrt 5

OpenStudy (anonymous):

x^2 = 5 x = sqrt 5

OpenStudy (anonymous):

\[3x^2=15\] You want to get divide 15 by 3 to get rid of the coefficient on the left side So: \[x^2=5\] Then you find the square root of that to get rid of the squared Your answer should be: \[x= \pm \sqrt{5}\]

OpenStudy (anonymous):

Okay, so then how would I solve 5x^2-3x=-8?

OpenStudy (anonymous):

add 8 to both sides

OpenStudy (anonymous):

then factor

OpenStudy (anonymous):

you might have to apply the quadratic formula here

OpenStudy (anonymous):

no, you don't. 5x^2 -5x + 8x + 8 = 0 i think you got the equation wrong, it should be 5x^2 - 3x = 8.

OpenStudy (anonymous):

oh, wait, you need the quadratic eq if it's 5x^2 - 3x = -8.

OpenStudy (anonymous):

oh really